OpenStudy (anonymous):

At the start of the basketball game, the referee tosses a ball for a jump ball. The equation that models the height of the ball is h(t) = -16t2 + 24t + 5 During what time interval is the height of the ball above 9 feet? (Round your answer to the nearest tenth of a second.) a) .4 seconds to 1.9 seconds b) .2 seconds to 1.3 seconds c) .6 seconds to 1.8 seconds d) .5 seconds to 2.1 seconds

OpenStudy (anonymous):

-16t2 + 24t + 5 = 9 -16t2 + 24t + 5 - 9 = 0 -16t2 + 24t - 4 = 0 using quad formula to solve for x its B